OpenStudy (anonymous):

If x^2=25-y^2, what is the value of the second derivative at the point (3, 4)?

OpenStudy (shubhamsrg):

did you find dy/dx first ?

OpenStudy (anonymous):

yeah, I got that as -x/y

OpenStudy (shubhamsrg):

seems correct, now dy/dx= -x/y differentiate again both sides with respect to x..you will get a dy/dx in the RHS..substitute (-x/y) in its place, and voila, 99% of the question is complete! ;)

OpenStudy (anonymous):

I'm confused, doesn't differentiating in the right get -x/dy?

OpenStudy (shubhamsrg):

ofcorse not,,how do you get that,, d/dx(-x/y) => use product rule or quotient rule of differentiation..

OpenStudy (shubhamsrg):

still stuck ?

OpenStudy (anonymous):

I think I'm closer. Using the quotient rule I got (x-dy-y)/y^2

OpenStudy (shubhamsrg):

seems you applied that wrong,,allow me to elaborate d/dx(-x/y) = [ y*d(-x)/dx - (-x)*(dy/dx) ] /y^2 is that clear so far ?

OpenStudy (anonymous):

yeah, that's what I got

OpenStudy (shubhamsrg):

so now you have d2y/dx2 = [ y*d(-x)/dx - (-x)*(dy/dx) ] /y^2 just simplify RHS,,use dy/dx= -x/y finally when all is done , substitute x=3 and y=4.. hope this helps..

OpenStudy (beginnersmind):

This can actually get a little messy. First, I'd use y' for dy/dx. Differentiating once you get 2x = 2yy' Don't express y' yet, just differentiate again. This gives 0 = 2(y')^2 + 2y'' or y'' = - (y')^2 You don't actually need the whole function y'' just its value at (3,4). So use the first equation to calculate y' at (3,4), then plug it in as a number. This is easier than getting y'' as a function of x and y.

OpenStudy (shubhamsrg):

yep..seconded!